OpenStudy (anonymous):

what is the square root of 50000

OpenStudy (perl):

its easier to first factor 5000 into prime factors

OpenStudy (anonymous):

what times what equals 50000

OpenStudy (anonymous):

What times it self is 50000?

OpenStudy (anonymous):

223.60679775

OpenStudy (theedwardsfamily):

223.60679775

OpenStudy (anonymous):

223.606798

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

misty i need help

OpenStudy (misty1212):

what question?

OpenStudy (perl):

hint: sqrt(50000) = sqrt( 5 * 10000 ) = sqrt(5) * sqrt(10000)

OpenStudy (anonymous):

223.606